OpenStudy (anonymous):

An audience of 450 people is seated in an auditorium. Each row contains the same number of seats and each seat in the auditorium is occupied. With three fewer seats per row, and five extra rows, the same audience could still be seated, occupying all seats. How many rows does the auditorium have?

OpenStudy (anonymous):

Sorry I tried and failed... :P

OpenStudy (anonymous):

oh wow, i got it by pure luck guessing. there are originally 18 seats per row and 25 rows. 18*25=450 (18-3)(25+5) =15*30 =450 as for the algebraic way to solve it, i'm not quite sure.

hero (hero):

To do it algebraically: xy = 450 (x-3)(y+5) = 450

hero (hero):

I can provide steps if you like

hero (hero):

I wonder how long it took loophole to figure it out by guessing

OpenStudy (anonymous):

second guess :P

hero (hero):

Good job either way